OpenStudy (anonymous):

Solve. 3x2 + 12x = 3

OpenStudy (anonymous):

Well first, put the equation in standard form -- ax^2 + bx + c = 0. Can you do that step?

OpenStudy (anonymous):

subtract three from both sides to get the equation equal to zero and then solve the quadratic equation like this: http://www.purplemath.com/modules/solvquad.htm

OpenStudy (anonymous):

no

OpenStudy (anonymous):

start with \[x^2+4x=1\] then \[(x+2)^2=1+4=5\] so \[x+2=\pm\sqrt{5}\] and finally \[x=-2\pm\sqrt{5}\]

OpenStudy (mertsj):

Didn't we solve this problem for you earlier this evening?

OpenStudy (anonymous):

no it gave me steps on how to do it then i got the wrong answer